### v1

@@ 3.3 Luminance Consistency Regularization
As our proposed hourglass covers an area of continuous unseen views, we conceptualize it as a bundle of multiple flipped diffuse reflection rays originating from the Lambertian surface (Fig. B in supplementary material). Furthermore, leveraging the property of Lambertian surfaces, which maintain constant luminance at a pixel regardless of the viewing angle, we propose a luminance consistency regularization.
@@
In addition to the existing outputs, our HourglassNeRF estimates the luminance $y$ as additional outputs per sample along a ray and renders the final luminance $\hat{y}$ by volume rendering as follows:
@@
The luminance map estimated from the hourglass is also optimized using the same $y_{\text{GT}}$ by minimizing the corresponding MSE loss $\tilde{l}_{\text{lum.}}$, resulting in our proposed training loss $\mathcal{L}_{\text{lum.}}=\eta l_{\text{lum.}}+\tilde{\eta}\tilde{l}_{\text{lum.}}$, where $\eta$ and $\tilde{\eta}$ are balancing weights.
By optimizing the volumetric rendered luminance, we can provide extra supervisory signals, whose GT luminances are obtained from GT pixel values without any burdensome process, to the blending weights $w$, which are directly related to the estimated depth map of NeRF. As a result, it leads to performance improvement under the few-shot scenario, and also enables a more physically consistent framework than FlipNeRF, adhering to the Lambertian surface assumption. Further details about the total loss and architecture are provided in the supplementary material.
@@ 4.1 Experimental Details
Datasets and metrics.
We evaluate our HourglassNeRF and other baselines on three representative datasets for novel view synthesis: Realistic Synthetic ${}^{\circ}$ [24], DTU [11], and Shiny Blender [40]. Realistic Synthetic ${}^{\circ}$ has 8 synthetic scenes, each containing 400 multi-view images with a white background. We use 4 and 8 views to train our HourglassNeRF, and conduct an analysis of the frequency regularization effect and the effect of luminance estimation within the 4-view setting, using the first 4 and 8 images from the training set for a fair comparison, following the protocol in [34, 33]. On the DTU dataset, featuring multi-view images with objects against a white table and black background, we compare HourglassNeRF against other methods across 3/6/9-view scenarios and perform an ablation study in the 3-view context, following [46]. Furthermore, we evaluate our method on Shiny Blender, which comprises 6 synthetic glossy objects, i.e. objects that are largely non-Lambertian.
@@
Baselines.
We compare our HourglassNeRF against the state-of-the-art (SOTA) regularization methods [9, 14, 25, 34, 45, 33] and the original mip-NeRF [1] on Realistic Synthetic ${}^{\circ}$. We also evaluate against pre-training [4, 5, 46] and regularization methods on the DTU dataset. On Shiny Blender, our HourglassNeRF is compared with FlipNeRF [33] and FreeNeRF [45], designed for ray augmentation and high-frequency regularization, respectively, as well as Ref-NeRF [40] tackling the view-dependent effects of non-Lambertian surfaces. The pre-training methods utilize DTU for pre-training, whereas regularization approaches and mip-NeRF are optimized per scene. Note that we report the results of other methods from [25, 34, 33], which outperformed the results from the corresponding original paper by modified training curriculum [25] and used the same training views to ensure a fair comparison [34, 33]. Kindly refer to our supplementary material for more experimental details.
4.2 Analysis of HourglassNeRF
@@
Figure 4: Comparison of fine details against FreeNeRF over the training phase. Compared to FreeNeRF which forcibly masks the high-frequency spectrum in the early training phase, our HourglassNeRF adaptively regularizes the high-frequency components of additional ray samples based on the target pixel photo-consistency (i.e. the angle between the original ray and hourglass) during the whole training process. As a result, our HourglassNeRF already achieves sharper fine details at 25K iteration than the fully trained FreeNeRF.
Figure 4: Comparison of fine details against FreeNeRF over the training phase. Compared to FreeNeRF which forcibly masks the high-frequency spectrum in the early training phase, our HourglassNeRF adaptively regularizes the high-frequency components of additional ray samples based on the target pixel photo-consistency (i.e. the angle between the original ray and hourglass) during the whole training process. As a result, our HourglassNeRF already achieves sharper fine details at 25K iteration than the fully trained FreeNeRF.
Figure 5: Comparison of our HourglassNeRF against multicasting strategy on DTU 3-view. Our HourglassNeRF outperforms FlipNeRF in all scenarios by a large margin, showing the superior efficiency as a ray augmentation method. The training time per scene is measured using the same GPU, iterations, and batch size. Note that FlipNeRF ($\psi=90^{\circ}$) experiences training instability when $\kappa>1$ due to an excessive number of additional training rays, which are inaccurately estimated during the initial training phase. The size of circles is proportional to $\kappa$, i.e. the number of augmented rays per original ray.
Frequency regularization effect of hourglass.
@@
Fig. 5 demonstrates the comparison between our HourglassNeRF and FreeNeRF, which is specialized for frequency regularization. Our HourglassNeRF renders sharper fine details than FreeNeRF from the earlier training phase and already achieves competitive rendering quality only after 25K training iterations compared to the fully trained FreeNeRF. Although FreeNeRF prevents overfitting by forcibly masking most of high-frequency components during early training, it results in non-satisfactory rendering outcomes that fail to capture fine details in the end. In contrast, our HourglassNeRF adaptively regularizes the high-frequency components of the hourglass samples based on the target pixel photo-consistency, preventing overfitting and enabling sharper rendering.
Effectiveness of hourglass as a bundle of rays.
In Fig. 5, we demonstrate the effectiveness of hourglass, which covers multiple unseen views, compared to FlipNeRF with a different number of augmented rays, i.e. the multicasting strategy. Based on FlipNeRF, we cast $\kappa$ extra rays per original ray $\mathbf{r}$, consisting of flipped reflection ray $\mathbf{r}^{\prime}$ and additional rays which are evenly spaced between $\mathbf{r}$ and its corresponding $\mathbf{r}^{\prime}$, i.e. within the area of our hourglass, toward the same target pixel. Note that we set the masking threshold $\psi=45^{\circ}$ identically for a fair comparison [^1] By using multicasting strategy with $\psi=90^{\circ}$, which is original masking threshold for FlipNeRF, the FlipNeRF suffers from severe training instability and fails to achieve comparable results. We conjecture that using an excessive number of augmented rays at the initial training phase, when the model is not sufficiently trained to generate valid additional rays, leads to the failure.. When $\kappa=1$, which amounts to the vanilla FlipNeRF, our HourglassNeRF outperforms FlipNeRF by a large margin, showing the efficacy of hourglass. However, exploiting $\kappa>1$ augmented rays, fails to achieve meaningful performance improvement. This suggests that our single hourglass casting, equipped with adaptive frequency regularization, is more effective than multicasting strategy over the same unseen area. Moreover, for a training efficiency, simply increasing the number of augmented rays in the multicasting strategy leads to longer training time due to the additional ray processing. However, our HourglassNeRF maintains training efficiency while improving performance by featurizing continuous unseen views with a single hourglass.
RegNeRF [25] FlipNeRF [33] FreeNeRF [45] HourglassNeRF ${}^{\dagger}$ Ground Truth
@@
Figure 6: Qualitative results of relative luminance maps. Our HourglassNeRF renders clearer relative luminance map than other baselines. With our explicitly estimated luminance $\hat{y}$, ${}^{\dagger}$ captures the detailed luminance on the handle more accurately. The red insets are visualized with doubled contrast for ease of comparison.
Table 2: Ablation study. Thanks to the proposed hourglass and $\mathcal{L}_{\text{lum.}}$, our HourglassNeRF achieves superior performance to its baseline FlipNeRF across all metrics.
@@
HourglassNeRF 16.87 0.886 0.074 0.088
${}^{\dagger}$ 17.18 0.891 0.062 0.079
Hourglass $\mathcal{L}_{\text{lum.}}$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Avg. $\downarrow$
@@
Table 2: Ablation study. Thanks to the proposed hourglass and $\mathcal{L}_{\text{lum.}}$, our HourglassNeRF achieves superior performance to its baseline FlipNeRF across all metrics.
Luminance estimation as an auxiliary task.
We show the quantitative and qualitative results of our proposed luminance estimation in Tab. 2 and Fig. 6, respectively. Since other baselines do not estimate the relative luminance $\hat{y}$ explicitly, their estimated luminance maps $\hat{y}^{\mathbf{c}}$ are derived from the estimated colors using Eq. 9 and we report our results of both $\hat{y}^{\mathbf{c}}$ and $\hat{y}$ for a fair comparison. Ours outperforms other representative methods based on the ray augmentation and frequency regularization. As illustrated in Fig. 6, our HourglassNeRF renders much clearer relative luminance maps than other baselines. Since the resulting relative luminances and pixel values share the identical blending weights $w$ for volumetric rendering, we are able to provide extra supervisory signals to the blending weights and effectively regularize them by the luminance estimation task.
4.3 Ablation Study
Tab. 2 shows the ablation study of our HourglassNeRF. By replacing the flipped reflection rays of FlipNeRF with our hourglass (1), it achieves performance improvement across most of the metrics, especially on SSIM and LPIPS by a large margin. However, it achieves rather degenerate results with only $\mathcal{L}_{\text{lum.}}$ (2). Since the flipped reflection rays are technically based on the specular reflection, our $\mathcal{L}_{\text{lum.}}$ goes against the physical concept of FlipNeRF training framework, leading to the performance drop. When we optimize our HourglassNeRF with our proposed $\mathcal{L}_{\text{lum.}}$ together ((1) $\rightarrow$ (3)), we can further enhance the rendering quality with clear fine details, especially on PSNR metric. Additional experiment of the masking threshold for hourglass is provided in our supplementary material.
4.4 Comparison with other Baselines
Realistic Synthetic ${}^{\circ}$.
As shown in Tab. 3 and Fig. 7, our HourglassNeRF achieves the SOTA performance over all the scenarios and metrics. Compared to RegNeRF and FlipNeRF, which utilize ray augmentation for training, ours renders superior quality of objects, with better capturing fine details. Even compared to FreeNeRF, ours shows a clearer texture of target objects, since high-frequency is adaptively regularized, not forcibly masked, during the early training.
Table 3: Quantitative results on Realistic Synthetic ${}^{\circ}$. Our HourglassNeRF outperforms other methods across all the scenarios and metrics.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average Error $\downarrow$
@@ 4.4 Comparison with other Baselines
HourglassNeRF 16.86 20.29 0.873 0.910 0.084 0.052 0.091 0.057
@@
Table 4: Quantitative results on DTU. Our HourglassNeRF outperforms its baseline FlipNeRF, and generally achieves competitive results. Kindly refer to Fig. 9 and our supplementary material for additional comparison against FreeNeRF without the W&B prior and pre-training methods, respectively. $\ddagger$ indicates the W&B prior.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average Error $\downarrow$

### v2

@@ 3.3 Luminance Consistency Regularization
We propose the luminance map as an effective additional training resource for few-shot scenarios with limited data, providing ’free lunch’ information easily derived from RGB images, and introduce luminance consistency regularization.
@@
In addition to the existing outputs, our ARC-NeRF estimates the luminance $y$ as additional outputs per sample along a ray and renders the final luminance $\hat{y}$ by volume rendering as follows:
@@
The luminance map estimated from the Area Ray is also optimized using the same $y_{\text{GT}}$ by minimizing the corresponding MSE loss $\tilde{l}_{\text{lum.}}$, resulting in our proposed training loss $\mathcal{L}_{\text{lum.}}=\eta l_{\text{lum.}}+\tilde{\eta}\tilde{l}_{\text{lum.}}$, where $\eta$ and $\tilde{\eta}$ are balancing weights.
By optimizing the volumetric rendered luminance, we can provide extra supervisory signals, whose GT luminances are obtained from GT pixel values without any burdensome process, to the blending weights $w$, which are directly related to the estimated depth map of NeRF. As a result, it leads to performance improvement under the few-shot scenario with more accurate object textures. Further details about the total loss and architecture are provided in the supplementary material.
@@ 4.1 Experimental Details
Datasets and metrics.   We evaluate our ARC-NeRF and other baselines on three representative datasets for novel view synthesis: Realistic Synthetic 360 [25], DTU [11], and Shiny Blender [42], focusing on the ability to capture fine details of objects. Realistic Synthetic 360 has 8 synthetic scenes, each containing 400 multi-view images with a white background. We use 4 and 8 views to train our ARC-NeRF, and conduct an analysis of the frequency regularization effect and the effect of luminance estimation within the 4-view setting, using the first 4 and 8 images from the training set for a fair comparison, following the protocol in [35, 34]. On the DTU dataset, featuring multi-view images with objects against a white table and black background, we compare ARC-NeRF against other methods across 3/6/9-view scenarios and perform an ablation study in the 3-view context, following [51]. Furthermore, we evaluate our method on Shiny Blender, which comprises 6 synthetic glossy objects.
@@
Figure 4: Comparison of fine details against FreeNeRF over the training phase. Compared to FreeNeRF which forcibly masks the high-frequency spectrum in the early training phase, ours adaptively regularizes the high-frequency components of additional ray samples based on the target pixel photo-consistency (i.e. the angle between the original ray and Area Ray) during the whole training process. As a result, our ARC-NeRF already achieves sharper fine details at 25K iteration than the fully trained FreeNeRF.
Baselines.   We compare our ARC-NeRF against the state-of-the-art (SOTA) regularization methods [9, 15, 26, 35, 50, 34] and the original mip-NeRF [1] on Realistic Synthetic 360. We also evaluate against pre-training [4, 5, 51] and more regularization methods [36, 47, 43, 48], including 3DGS-based approaches [14, 48] which have become another mainstream, on the DTU dataset. On Shiny Blender, our ARC-NeRF is compared with FlipNeRF [34] and FreeNeRF [50], designed for ray augmentation and high-frequency regularization, respectively, as well as Ref-NeRF [42] addressing the ray parameterization for view-dependent effects. The pre-training methods utilize DTU for pre-training, whereas regularization approaches and mip-NeRF are optimized per scene. Note that we report the results of other methods from [26, 35, 34], which outperformed the results from the corresponding original paper by modified training curriculum [26] and used the same training views to ensure a fair comparison [35, 34]. Kindly refer to our supplementary material for more details.
4.2 Analysis of ARC-NeRF
@@
Frequency regularization effect of Area Ray.   Fig. 4 demonstrates the comparison between our ARC-NeRF and FreeNeRF, which is specialized for frequency regularization. Ours renders sharper fine details than FreeNeRF from the earlier training phase and already achieves competitive rendering quality only after 25K training iterations compared to the fully trained FreeNeRF. Although FreeNeRF prevents overfitting by forcibly masking most of high-frequency components during early training, it results in non-satisfactory rendering outcomes that fail to capture fine details in the end. In contrast, ours adaptively regularizes the high-frequency components of the Area Ray samples based on the target pixel photo-consistency, preventing overfitting and enabling sharper rendering.
Figure 5: Comparison of our ARC-NeRF against multicasting strategy on DTU 3-view. Our ARC-NeRF outperforms FlipNeRF in all scenarios by a large margin. The training time per scene is measured using the same GPU, iterations, and batch size. The size of circles is proportional to $\kappa$, i.e. the number of augmented rays per original ray.
RegNeRF [26] FlipNeRF [34] FreeNeRF [50] ARC-NeRF ARC-NeRF Ground Truth
@@
Figure 6: Qualitative results of relative luminance maps. Our ARC-NeRF renders clearer relative luminance map than other baselines. With our explicitly estimated luminance $\hat{y}$, ours captures the detailed luminance on the handle more accurately. The red insets are visualized with doubled contrast for ease of comparison.
Effectiveness of Area Ray as a bundle of rays.   In Fig. 5, we demonstrate the effectiveness of Area Ray, which covers a broad range of unseen views, compared to FlipNeRF with a different number of augmented rays, i.e. the multicasting strategy. Based on FlipNeRF, we cast $\kappa$ extra rays per original ray $\mathbf{r}$, consisting of flipped reflection ray $\mathbf{r}^{\prime}$ and additional rays which are evenly spaced between $\mathbf{r}$ and its corresponding $\mathbf{r}^{\prime}$, i.e. within the area of our Area Ray, toward the same target pixel. When $\kappa=1$, which amounts to the vanilla FlipNeRF, our ARC-NeRF outperforms FlipNeRF by a large margin, showing the efficacy of Area Ray. However, exploiting $\kappa>1$ augmented rays, fails to achieve meaningful performance improvement. This suggests that our single Area Ray casting, equipped with adaptive frequency regularization, is more effective than multicasting strategy over the same unseen area. Moreover, for a training efficiency, simply increasing the number of augmented rays in the multicasting strategy leads to longer training time due to the additional ray processing. However, our ARC-NeRF maintains training efficiency while improving performance by featurizing continuous unseen views with a single Area Ray. Note that we set the masking threshold $\psi=45^{\circ}$ identically for a fair comparison [^1] By using multicasting strategy with $\psi=90^{\circ}$, which is original masking threshold for FlipNeRF, the FlipNeRF suffers from severe training instability and fails to achieve comparable results. We conjecture that using an excessive number of augmented rays at the initial training phase, when the model is not sufficiently trained to generate valid additional rays, leads to the failure..
@@
ARC-NeRF 16.87 0.886 0.074 0.088
ARC-NeRF 17.18 0.891 0.062 0.079
Luminance estimation as an auxiliary task.   We show the quantitative and qualitative results of our proposed luminance estimation in Tab. 1 and Fig. 6, respectively. Since other baselines do not estimate the relative luminance $\hat{y}$ explicitly, their estimated luminance maps $\hat{y}^{\mathbf{c}}$ are derived from the estimated colors using Eq. 9 and we report our results of both $\hat{y}^{\mathbf{c}}$ and $\hat{y}$ for a fair comparison. Ours outperforms other representative methods based on the ray augmentation and frequency regularization. As illustrated in Fig. 6, our ARC-NeRF renders much clearer relative luminance maps than other baselines. Since the resulting relative luminances and pixel values share the identical blending weights $w$ for volumetric rendering, we are able to provide extra supervisory signals to the blending weights and effectively regularize them by the luminance estimation task.
Ablation study.   Tab. 2 shows the ablation study of our ARC-NeRF. By replacing the flipped reflection rays of FlipNeRF with our Area Ray (1), it achieves performance improvement across most of the metrics, especially on SSIM and LPIPS by a large margin. However, it achieves rather degenerate results with only $\mathcal{L}_{\text{lum.}}$ (2). Since the flipped reflection rays are technically based on the mirror reflection, which assumes that the pixel’s luminance varies depending on its viewpoint, our $\mathcal{L}_{\text{lum.}}$ goes against the physical concept of FlipNeRF training framework, leading to the performance drop. When we optimize our ARC-NeRF with our proposed $\mathcal{L}_{\text{lum.}}$ together ((1) $\rightarrow$ (3)), we can further enhance the rendering quality with clear fine details, especially on PSNR metric. Additional experiment of the masking threshold for Area Ray is provided in our supplementary material.
Area Ray $\mathcal{L}_{\text{lum.}}$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Avg. $\downarrow$
@@
Table 2: Ablation study. Thanks to the proposed Area Ray and $\mathcal{L}_{\text{lum.}}$, our ARC-NeRF achieves superior performance to its baseline FlipNeRF across all metrics.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Avg. $\downarrow$
@@
ARC-NeRF 16.86 20.29 0.873 0.910 0.084 0.052 0.091 0.057
Table 3: Quantitative results on Realistic Synthetic 360. ARC-NeRF outperforms others across all the scenarios and metrics.
4.3 Comparison with other Baselines
Realistic Synthetic 360.   As shown in Tab. 3 and Fig. 7, our ARC-NeRF achieves the SOTA performance over all the scenarios and metrics. Compared to RegNeRF and FlipNeRF, which utilize ray augmentation, ours renders superior quality of objects, with better capturing fine details. Even compared to FreeNeRF, ours shows a clearer texture of target objects, since high-frequency is adaptively regularized, not forcibly masked, during the early training.
mip-NeRF [1] MixNeRF [35] FreeNeRF [50] FlipNeRF [34] ARC-NeRF Ground Truth
(a) Realistic Synthetic 360 4-view
mip-NeRF [1] RegNeRF [26] FreeNeRF [50] FlipNeRF [34] ARC-NeRF Ground Truth
(b) DTU 3-view
Ref-NeRF [42] FreeNeRF [50] FlipNeRF [34] ARC-NeRF Ground Truth
(c) Shiny Blender 4-view
Figure 7: Qualitative comparisons on Realistic Synthetic 360, DTU, and Shiny Blender. Our ARC-NeRF achieves notable rendering quality with better capturing fine details. $\ddagger$ indicates the W&B prior. Additional results are provided in the supplementary material.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Avg. $\downarrow$
Mip-NeRF [1] 8.68 0.571 0.353 0.323
3DGS [14] 14.18 0.628 0.301 0.191
Pre-training.
PixelNeRF [51] 16.82 0.695 0.270 0.147
PixelNeRF [51] 18.95 0.710 0.269 0.125
SRF [5] 15.32 0.671 0.304 0.171
SRF [5] 15.68 0.698 0.281 0.162
MVSNeRF [4] 18.63 0.769 0.197 0.113
MVSNeRF [4] 18.54 0.769 0.197 0.113
Regularization.
DietNeRF [9] 11.85 0.633 0.314 0.243
RegNeRF [26] 18.89 0.745 0.190 0.112
MixNeRF [35] 18.95 0.744 0.203 0.113
SimpleNeRF [36] 16.25 0.751 0.249 0.143
DiffusioNeRF [47] 16.20 0.698 0.160 0.128
SparseNeRF [43] 19.55 0.769 0.201 0.102
FreeNeRF [50] 19.92 0.781 0.125 0.086
FreeNeRF [50] 19.23 0.769 0.149 0.103
FlipNeRF [34] 19.55 0.767 0.180 0.101
SparseGS [48] 18.89 0.702 0.229 0.117
ARC-NeRF 19.85 0.773 0.146 0.096
@@ 4.3 Comparison with other Baselines
Table 4: Quantitative results on DTU 3-view. Our ARC-NeRF outperforms its baseline FlipNeRF, and generally achieves competitive results. Kindly refer to our supplementary material for additional comparison under 6/9-view scenarios. $\dagger$ and $\ddagger$ indicate the fine-tuning and the W&B prior, respectively.
DTU.   Tab. 4 shows the quantitative comparisons on DTU. Our ARC-NeRF outperforms the baseline FlipNeRF, and achieves competitive results overall. As demonstrated in Fig. 7, ours shows notable outcomes with sharply rendered textures. FreeNeRF is trained with the black and white prior assuming the estimated black and white color as the background and table, respectively, which is a highly strong assumption specific to the dataset, and achieves degenerate results without the prior. However, our ARC-NeRF achieves competitive performance without any heuristic prior by using Area Ray, which enables adaptive regularization of high-frequency. Kindly refer to the supplementary material for its qualitative comparison.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average $\downarrow$
Ref-NeRF [42] 17.10 0.821 0.190 0.142
FreeNeRF [50] 16.99 0.828 0.157 0.131
FlipNeRF [34] 18.14 0.847 0.141 0.109
ARC-NeRF 18.68 0.851 0.141 0.107
Table 5: Quantitative results on Shiny Blender 4-view. While FlipNeRF demonstrates competitive performance by utilizing flipped mirror reflection rays, which are physically aligned with the non-Lambertian surfaces, our approach consistently outperforms FlipNeRF and other comparative methods.
Shiny Blender.   We additionally compare our ARC-NeRF and others on Shiny Blender, which is largely non-Lambertian, as in Fig. 7(c) and Tab. 5. Although FlipNeRF achieves competitive performance thanks to its ray augmentation utilizing mirror reflection, which accords with the non-Lambertian surfaces, ours still outperforms others including FlipNeRF. We conjecture that under the few-shot setting where the major challenge is to learn 3D geometry effectively while preventing the overfitting, our adaptive high-frequency regularization plays an important role.
5 Conclusion
In this work, we have approached the few-shot novel view synthesis from the perspective of a ray parameterization. Our ARC-NeRF casts an Area Ray as an augmented ray, which effectively prevents overfitting by adaptively regularizing the high-frequency components based on the target pixel photo-consistency. Furthermore, as a single Area Ray covers a broader range of unseen view areas than an original ray, our ARC-NeRF is trained more efficiently without casting multiple augmented rays per original ray, maximizing the use of ray augmentation. We also utilize a free lunch training resource, i.e. luminance map, to provide additional supervision for output densities, further improving the fine details. Our ARC-NeRF achieves the SOTA or competitive performance compared to other methods under various few-shot scenarios. We expect that our ARC-NeRF is able to inspire a meaningful approach in the few-shot NeRF.
Acknowledgements.   This work was supported by NRF grant (2021R1A2C3006659) and IITP grant (RS-2021-II211343), both funded by MSIT of the Korean Government. The work was also supported by Samsung Electronics (IO201223-08260-01).
Supplementary Material
Hyperparameter Realistic Synthetic 360 [25] DTU [11] Shiny Blender [42] & Balancing Weights 4-view 8-view 3-view 6-view 9-view 4-view LR $[1\mathrm{e}{-3},1\mathrm{e}{-5}]$ $[2\mathrm{e}{-3},2\mathrm{e}{-5}]$ $[1\mathrm{e}{-3},1\mathrm{e}{-5}]$ Warm-up Iter. 512 1024 512 2048 512 $\eta_{\text{Ori.}}(\text{for }\mathcal{L}_{\text{Ori.}})$ $1\mathrm{e}{-1}$ $1\mathrm{e}{-2}$ $1\mathrm{e}{-1}$ $1\mathrm{e}{-2}$ $1\mathrm{e}{-1}$ $\eta_{\text{lum.}}(\text{for }l_{\text{lum.}})$ $1\mathrm{e}{-3}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-3}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-3}$ $\tilde{\eta}_{\text{lum.}}(\text{for }\tilde{l}_{\text{lum.}})$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-6}$ $1\mathrm{e}{-4}$
Table A: Hyperparameters and balancing weights. Since our ARC-NeRF is built upon FlipNeRF, we follow the training details for other hyperparameters, which are not mentioned here, as FlipNeRF. $[\alpha,\beta]$ denotes the annealing from $\alpha$ to $\beta$.
A Experimental Setting
Figure A: Network architecture of ARC-NeRF. Our ARC-NeRF estimates the additional output $y$, i.e. the relative luminance for our auxiliary luminance estimation task.
Implementational details.
Our ARC-NeRF is implemented upon FlipNeRF [34], and we follow its overall training scheme. We utilize the scene space annealing strategy during the initial training phase following [26, 35, 34]. Furthermore, we adopt the initial warm up and exponential decay for the learning rate. We use the Adam optimizer [16] with gradient clipping set to 0.1 for both each element of the gradient value and the gradient’s norm. Our ARC-NeRF is trained for 500 pixel epochs using a batch size of 4,096 on four NVIDIA RTX 3090 GPUs. Additionally, since our proposed Area Ray encompasses broader areas of unseen views compared to a single ray, we set the masking threshold $\psi$ as 45, which is smaller than that of FlipNeRF, to avoid over-regularization effect of augmented rays. The related experiment is demonstrated in Tab. C.
Hyperparameters.
For additional details on hyperparameters and loss balancing terms based on training views and datasets, kindly refer to Tab. A. Note that our ARC-NeRF follows the same hyperparameters as FlipNeRF for other training losses and schemes which are not specified in Tab. A.
B Further Details of Method
Architectural details.
Our ARC-NeRF leverages the network architecture of mip-NeRF [1], which is commonly used in several few-shot NeRF models [50, 26, 34, 35]. Moreover, our ARC-NeRF additionally estimates the relative luminance $y$. Kindly refer to more details in Fig. A.
Total loss.
Our ARC-NeRF is trained to maximize the log-likelihood of the target pixel $\mathbf{c}_{\text{GT}}$ for both sets of original input rays $\mathcal{R}$ and our proposed Area Rays $\tilde{\mathcal{R}}$, as well as to minimize the mean squared errors (MSE) between the ground-truth and estimated pixel values. Except our proposed $\mathcal{L}_{\text{lum.}}$, we use the same training losses as those of FlipNeRF. Note that we use $\mathcal{L}_{\text{MSE}}$ only for $\mathcal{R}$ and exploit a batch of Area Rays instead of flipped reflection rays. Summing up, the total loss over a batch is calculated as follows:
$\begin{split}\mathcal{L}_{\text{Total}}&=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{lum.}}+\eta_{\text{NLL}}\mathcal{L}_{\text{NLL}}+\tilde{\eta}_{\text{NLL}}\tilde{\mathcal{L}}_{\text{NLL}}\\ &+\eta_{\text{UE}}\mathcal{L}_{\text{UE}}+\tilde{\eta}_{\text{UE}}\tilde{\mathcal{L}}_{\text{UE}}^{\prime}+\eta_{\text{BFC}}\mathcal{L}_{\text{BFC}}+\eta_{\text{Ori.}}\mathcal{L}_{\text{Ori.}},\\ &\text{where}\quad\mathcal{L}_{\text{lum.}}=\eta_{\text{lum.}}l_{\text{lum.}}+\tilde{\eta}_{\text{lum.}}\tilde{l}_{\text{lum.}}.\end{split}$ (12)
$\eta$’s and $\tilde{\eta}$’s represent the loss balancing weights for the original input rays and additional Area Rays, respectively.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average Err. $\downarrow$
FlipNeRF [34] 16.47 0.866 0.091 0.095
ARC-NeRF
   w/o view. jitter. 16.66 0.869 0.087 0.093
   w/ view. jitter. 16.86 0.873 0.084 0.091
Table B: Effect of Viewing direction jittering. On Realistic Synthetic 360 4-view, we are able to achieve marginal performance improvement while still outperforming FlipNeRF without the jittering strategy.
$\psi$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average Err. $\downarrow$
180 (None) 18.15 0.749 0.179 0.120
90 18.63 0.762 0.163 0.110
75 19.02 0.764 0.156 0.105
60 18.94 0.766 0.154 0.105
45 19.85 0.773 0.146 0.096
30 18.78 0.765 0.160 0.107
15 18.65 0.761 0.163 0.111
[…]
C Additional Experiments
Viewing direction jittering.
For Realistic Synthetic 360 [25] and Shiny Blender [42], which consist of inward-facing synthetic scenes with objects located at the center, we adopt the viewing direction jittering, which is a minor additional strategy slightly improving the performance. We simply add the Gaussian random noise to the input viewing direction $\mathbf{d}$ to improve the robustness for the slight change of viewpoints. As shown in Tab. B, we are able to achieve marginal improvement of rendering quality while still outperforming its baseline, FlipNeRF, even without the jittering strategy.
Masking thresholds.
Our ARC-NeRF utilizes an additional batch of Area Rays covering a broader area of unseen views, and the high-frequency components of samples along an Area Ray are adaptively regularized via Integrated Positional Encoding (IPE) depending on the angle between the original input direction and the estimated normal vector, i.e. the target pixel photo-consistency. As a result, with the same $\psi=90^{\circ}$ as FlipNeRF, our ARC-NeRF might suffer from the performance degradation due to over-regularization. As demonstrated in Tab. C, our ARC-NeRF achieves the best result with $\psi=45^{\circ}$. The larger $\psi$ becomes than $45^{\circ}$, the worse the performance, as an Area Ray covering too wide area of unseen views leads to over-regularization, which adversely affects the training. On the other hand, a smaller $\psi$ than $45^{\circ}$ also leads to poorer performance, as the newly generated Area Rays are excessively filtered, resulting in only a limited number of augmented Area Rays being utilized for training. Note that the masking threshold $\psi$ depends on the characteristics of casting ray rather than being the hyperparameter which needs to be finetuned elaboratively.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average $\downarrow$
FreeNeRF [50] 19.92 0.781 0.125 0.086
FreeNeRF [50] 19.23 0.769 0.149 0.103
ARC-NeRF 19.85 0.773 0.146 0.096
FreeNeRF [50] FreeNeRF [50] ARC-NeRF
Figure B: Quantitative and qualitative comparison with FreeNeRF on DTU 3-view. Although FreeNeRF achieves high-quality of rendering with only a few images, it depends on the white and black prior, which is highly heuristic based on the characteristics of dataset. Our ARC-NeRF achieves comparable performance to FreeNeRF without the heuristic prior for training. $\ddagger$ denotes the W&B prior.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Avg. Err. $\downarrow$